Starting from rest, a prototype car accelerates to 50m/s in 12 seconds. Find its acceleration and distance it travels over the 12 seconds.

Formulas : x= x(0)+v(0)t + 1/2at^2

v=v(0)+at

a=?

x=?
Answer #1

Answer:

First, we will find the acceleration of car using v = v(0) + at

As the car starts from rest, v(0) = 0

Now, final velocity after 12 seconds is 50m/s

So, equation is : 50 = 0 + a*12

So, a = 50/12 = 4.166 m/s^2

Now, lets find the distance travelled by the car:

x = x(0) + v(0)t + 1/2 * (at^2)

as it starts from rest x(0) and v(0) are 0.

As, a= 50/12 and t = 12. So,

x = (1/2) * (50/12) * 12^2

x = (1/2) * (50/12) * 144

x = 300 metres.

PLEASE UPVOTE IF YOU FOUND THIS HELPFUL!